OpenStudy (anonymous):

A triangle has an area equal to 1/2(12x^2y^2)(6x^3y^5). Which of the following expressions represents the simplified form of the area of this triangle? Select one: A. 72x^2y^2 B. 36x^5y^7 C. 36x^6y^10 D. 36x3^y^6

OpenStudy (anonymous):

@hoblos could you help?

OpenStudy (hoblos):

multiply the numbers and add the powers of like variables so (1/2)*12*6 = 36 x^(2+3) y^(2+5) you get 36x^5y^7

OpenStudy (anonymous):

so you just add up the exponents?

OpenStudy (hoblos):

yes (x^a)(x^b) = x^(a+b)